on top of a spiral spring of force constant 509N is placed placed in mass of 5x10raise to power 3 . If the spring is Compressed down ward by alength of 0.02m and then release Calculate the height to which the mass projected 


1
Expert's answer
2022-03-02T14:38:41-0500

According to the energy conservation, the initial potential energy of the spring is equal to the final gravitational potential energy of the body:


"\dfrac{kx^2}{2} = mgh"

where "k = 509N\/m, x = 0.02m, g = 9.8m\/s^2, m=5000kg" and "h" is required height. Thus, obtain:


"h = \dfrac{kx^2}{2mg} = \dfrac{509\cdot 0.02^2}{2\cdot 5000\cdot 9.8} \approx 2\times 10^{-9}m"